OpenStudy (anonymous):

Identify the 31st term of an arithmetic sequence where a 1 = 26 and a22 = -226. -334 -274 -284 -346

OpenStudy (anonymous):

Similar to the earlier one first find d using a22 = a + 21d -226 = 26 + 21d -226 - 26 = 21d -252 = 21d d = -252/21 d = -12 a31 = a + 30d a31 = 26 + 30*-12 = 26 - 360 = -334 answer is -334